OpenStudy (anonymous):

Does anybody have an idea of what this means? Explain the steps you would follow to evaluate 40 + 4h =96 5 where h =3D 8.

OpenStudy (anonymous):

could you retype the questions please? I don't really understand the format either. Are you asking: 40 + 4h = 965 where h=24D?

OpenStudy (anonymous):

My teacher sent it through email :P S

OpenStudy (anonymous):

She probably made a mistake O.o

OpenStudy (anonymous):

T_T

OpenStudy (austinl):

I assume it is similar to \(f(h)=40+4h=965\) \(g(D)=24D\) \(f(g(D))=~?\) are these numbers correct?

OpenStudy (anonymous):

I just got another email from my teacher :D She said she is having a live session for these questions she sent me on Thursday :P

OpenStudy (anonymous):

It's okay guys :D Thanks!

OpenStudy (austinl):

Okay, well I would take full advantage of that session.

OpenStudy (anonymous):

Thank goodness I don't have to do these questions :P

OpenStudy (anonymous):

@austinL, what do you mean?

OpenStudy (austinl):

I was meaning, pay attention to what she says. This has me stumped, but likely from the format of the question.

OpenStudy (anonymous):

I know. Probably some mistake ;)

OpenStudy (anonymous):

Anyways, thanks guys :D